OpenStudy (anonymous):

f-1(x) =??? {x/2, x ≤ 0 {sqrt(x), x>0

OpenStudy (anonymous):

\[f(x) = \left\{\begin{matrix} x/2 & x \le 0\\ \sqrt{x} & x > 0 \end{matrix}\right.\] is this f(x) ?